22. The length of a rectangle is 2 in. more than twice its width. If the perimeter of the rectangle is 52 in., find the width of the rectangle.

Answer by checkley71(8403)

2(2W+2)+2W=52 OR 4W+4+2W=52 OR 6W=52-4 OR 6W=48 OR W=48/6 OR W=8IN FOR THE WIDTH.
PROOF 2(2*8+2)+2*8=52 OR 2(16+2)+16=52 OR 2(18)+16=52 OR 36+16=52 OR 52=52